OpenStudy (anonymous):

What is (a+5)/5a = 10, how do you solve for a?

OpenStudy (anonymous):

is this \[\frac{a+5}{5a}=10\]?

OpenStudy (anonymous):

yes, that is it...

OpenStudy (anonymous):

then multiply both side by 5a to get \[a+5=50a\]

OpenStudy (anonymous):

\[5=49a\] \[a=\frac{5}{49}\]

OpenStudy (anonymous):

ok, thanks. what about (z-4)/(z+4)= 2?